OpenStudy (anonymous):

Write the polynomial in standard forms, and identify the zeros of the function and the x-intercepts of its graph. f(x)=(x+2)(x- sqrt.3i)(x+sqrt.3i)

OpenStudy (anonymous):

\[f(x)=(x+2)(x-\sqrt{3i})(x+\sqrt{3i})\]

OpenStudy (anonymous):

\[(x-\sqrt{3}i)(x+\sqrt{3}i)=x^2+9\] is a start then multiply \[(x+2)(x^2+9)\] to get your answer in standard form

OpenStudy (anonymous):

Thanks, I already got that done. How would I identify the zeros?

OpenStudy (anonymous):

Oh, nevermind. I got it now.

OpenStudy (anonymous):

yeah, they are displayed for you